# Multimodal Fusion on Low-quality Data: A Comprehensive Survey

### v1

@@ 4.2 Quality-discrepancy methods
In addition to the issue of discrepancy in unimodal learning property, even though all modalities convey the same concepts, they contain varying amounts of information related to the target event or objects, resulting the disrepancy in unimodal data quality. Given the inherently greedy nature of deep neural networks, recent studies have further uncovered that multimodal models tend to heavily rely on a single modality with high quality that provides sufficient target-related information, while neglecting the others, leading to their under-optimized [8, 10]. As shown in Fig. 4, Peng et al. [8] empirically verified that the quality of unimodal representations in the multimodal model is worse than that in the corresponding unimodal models, despite the overall superior performance of the multimodal models. Moreover, this under-optimization of each modality is imbalanced, and the modality with high quality is less under-optimized. They claim that this imbalanced problem is caused by the model preference for the high-quality modality with better performance (i.e., the dominant modality). For instance, the audio modality in the curated sound-oriented dataset, VGGSound [130], is less under-optimized compared to the vision modality. Theoretical analysis conducted by Huang et al.  [10] has further confirmed that for the jointly trained multimodal late-fusion framework, the encoder networks can only learn a subset of modalities, and other modalities may fail to be discovered well. This finding is consistent with empirical observations, indicating that the imbalanced preference for different modalities issue needs to be addressed in multimodal learning. To tackle this problem brought by discrepancy in unimodal data quality, a series of studies have proposed, encompassing learning-objective-based method, optimization-based method, architecture-based method and data-argumentation-based method.
@@ 4.2.1 Learning-objective-based method
Considering the preference of multimodal model for modality with high quality, several methods which introduce additional learning objective beyond normal joint learning loss function are proposed. These methods tend to utilize additional loss to break the modality preference of multimodal model:
@@
where $\mathcal{L}_{mm}$ is the multimodal loss function and $\mathcal{L}_{uni}$ is specific designed loss function that improving unimodal learning, especially the learning of modality with lower quality. Concretely, Yang et al. [117] endowed the multimodal contrastive learning loss to constrain the audio and visual modality to be as close as possible in the semantic space, then help to make full use of all modalities. Also paying attention to the multimodal semantic space, Ma et al[131] proposed a novel regularization technique Calibrating Multimodal Learning, to enforces the consistency between prediction confidence and the number of modalities, avoiding the issue that multimodal classification models tend to over-rely on specific modalities. In addition to these methods in semantic space, some methods introduce the learning objective of knowledge distillation. Du et al. [118] proposed to distill the pre-trained uni-modal features to the corresponding parts of multimodal late-fusion models, to avoid the learning preference for one modality. Differently, Liu et al. [119] designed self-distillation training strategy, which can automatically teach less-optimized modality from better-optimized modality. Besides these common multimodal classification task, Xu et al [120] focused on the audio-visual fine-grained classification task, which have a higher demand for distinguishable feature distribution. They proposed the multimodal Cosine loss, MMCosine, which performs a modality-wise $L2$ normalization to features and weights towards balanced and better multimodal fine-grained learning. Beyond the generalization performance, Yang et al [121] put attention on the multi-modal robustness. They emphasized that multi-modal models are often vulnerable to attacks on the specific modality, and introduced margin-based regularization to ensure multi-modal robustness.
@@ 4.2.2 Optimization-based method
Optimization-based methods focus on the back-propagation stage. These methods respectively pay attention to the magnitude and direction of unimodal gradient, to improve the learning of modality with lower quality. To slow down the learning of preferred modality with high quality and accordingly offer more training efforts to others modalities, Peng et al. [8] proposed an on-the-fly gradient modulation strategy, which dynamically monitors the contribution difference of various modalities to the final prediction during training and mitigates the gradient magnitude of dominant modality to focus more on other modalities. Inspired by this gradient magnitude modulation idea, Sun et al. [123] extended it into the graph-based models of rumor detection task, adaptively adjusting the magnitude of gradients to alleviate the preference for high-quality modality. Similarily, Fu et al. [124] extended the idea of gradient magnitude modulation into audio-visual video parsing task. They further proposed a modality-separated decision unit to well evaluate the unimodal prediction in the more complex audio-visual video parsing framework with more cross-modal interaction. Besides the gradient magnitude, the direction of unimodal gradient is also explored. Fan et al. [122] suggested that the slow-learning modality’s updating direction is severely disturbed by the preferred high-quality modality, making it hard to its training. Subsequently, they utilize prototypes, which are the centroids of each class within the representation space, to refine the updating direction, aiming to improve unimodal performance.
@@ 4.2.3 Architecture-based method
Architecture based methods aim to balance the training of modalities with discrepancy in quality by designing better unimodal representation learning modules. He et al. [125] designed a Multimodal Temporal Attention module, which takes into account the temporal influences of all modalities on each individual unimodal branch. This module facilitates the interaction among unimodal branches and achieves adaptive inter-modal balance. Su et al. [126] focused on the generative scenarios, and utilized the coordinated feature space, named Coordinated Knowledge Mining, to improve the under-optimized modalities with the guidance of preferred high-quality modalities. Lin et al. [127] proposed a novel Variational Feature Fusion Module, which regards fusion features as random variables and obtains more balanced segmentation performance for different categories and modalities.
4.2.4 Data-argumentation-based method
Data-argumentation-based methods are with the idea of enhancing low-quality modality at the data input stage. The basic intuition is that solely training of unimodal data could avoid the influence of other modalities. There are two kind of strategies, introducing additional individual training for less learnt low-quality modality or removing the data of better learnt high-quality modality. For the former idea, Wu et al. [9] firstly measured the relative speed at which the model learns from one modality compared to other modalities, and then introduce the re-balancing steps at which update the underutilized uni-modal branches intentionally in order to accelerate the model to learn from its input modality. For the latter one, Zhou et al. [128] suggested to adaptively mask the data input of well learnt modality, accordingly encouraging the model to fit other modalities better. Most imbalanced multi-modal learning methods ignore the fine-grained modality discrepancy, where the dominant modality could vary across different samples under realistic scenarios. Wei et al. [129] introduced a Shapley-based sample-level modality valuation metric, to observe the contribution of each modality during prediction for each sample, and then dynamically re-sample low-contributing modalities during training. This method achieves fine-grained multi-modal cooperation.
@@ 4.3 Discussion
Overall, balanced multimodal learning methods focus on the discrepancy in learning property or data quality among different modalities, which is caused by the heterogeneity of multimodal data. These methods propose solutions from different perspectives, including learning objective, optimization, architecture and data argumentation. The balanced multimodal learning is still a sunrise field, with plenty of under-explored directions. For example, it is promising to explore the cooperation between modalities with theoretical guidance. Moreover, the current methods mainly limit to the typical multimodal task, mostly discriminative tasks and a few generative tasks. In fact, beyond these, the Multimodal Large Language Model, which also needs to jointly integrate different modalities, could also suffer from this imbalance problem. It is expected to extend current studies or design new solution in Multimodal Large Language Model settings.
@@ 5 Dynamic Multimodal Fusion
Current multimodal fusion methods often based on the assumption that multimodal data are of static quality, which does not always hold in real-world scenarios. Handling multimodal data with dynamically varying quality is an inevitable problem for multimodal intelligence systems. Due to the unexpected environmental factors and sensor issues, some modalities may suffer from poor reliability and loss task-specific information. Besides, the quality of different modalities are dynamically varying from scenarios as shown in Figure 5. This phenomenon motivates a new multimodal learning paradigm, i.e., dynamical multimodal fusion, which aims to adapt to the dynamically changing quality of multimodal data and selectively integrate task-specific information. In this section, we focus on the challenge on dynamical multimodal fusion and categorize current advances in this literature into three main lines including heuristic, attention-based and uncertainty-aware dynamic fusion.
@@ 5.1 Heuristic dynamic fusion
As a concrete example, illumination conditions can serve as a criterion to introduce dynamic fusion. In around-the-clock applications, previous work identifies that the quality of RGB and thermal modalities are temporally varying in multispectral pedestrian detection task [133]. While the RGB modality tends to contain more useful information compared to the thermal modality under normal illumination conditions, this relationship can be reversed under low light or nighttime conditions where the thermal modality becomes more reliable than RGB. Based on this observation, the authors proposed to utilize an illumination aware fusion module that fuse features from the two modalities adaptively. Similarly, to enable adaptation to varying illumination conditions, Guan et al. [132] proposed an illumination-aware weighting mechanism that evaluates the lighting of each input scene. These illumination weights are utilized to dynamically integrate the outputs of sub-networks specialized for day and night detection and segmentation. The proposed method can further be applied to pedestrian detection in security surveillance and autonomous driving systems.
@@
Besides environmental factors like illumination, the criterion of dynamic fusion can be devised according to knowledge about the properties of neural networks. When vision models are involved, activity level of the feature maps and the scaling factors in batch-norm layers can indicate input informativeness. Inspired by this, Li et al. [134] proposed an encoder-decoder style network with dynamic fusion strategy. In their implementation, the l1-norm of feature maps is calculated as channel-wise importance, which is then normalized to form a weighting vector that selectively emphasizes informative features from each modality. Wang et al. [112] proposed Channel Exchanging Networks (CEN) that dynamically exchange information between modality sub-networks based on channel importance evaluated by the scaling factors of batch-normalize layers. By performing channel exchange within modality-specific regions and sharing convolutional filters, CEN balances unimodal feature learning and multimodal fusion dynamically.
@@
Some relevant works also utilize task-specific modules to achieve dynamic fusion. To highlight a few, Panda et al. [135] introduces AdaMML, an adaptive multimodal learning in efficient video recognition. By dynamically selecting optimal input modalities per video segment, this work contributes to data-dependent multimodal selection for efficient video recognition. In this work, a lightweight policy network is devised to dynamically select optimal input modalities for each video segment. AdaMML utilizes Gumbel-Softmax sampling for differentiable training of discrete decisions. Similarly, Xue et al. [137] proposed DynMM, a novel approach for dynamic multimodal fusion in deep learning. It introduces modality-level and fusion-level decisions, allowing adaptive selection of input modalities and fusion operations based on data characteristics.
@@ 5.2 Attention-based dynamic fusion
Self-attention models dependency within the input sequence by allowing each element to interact with others. In multimodal emotion recognition (involving textual and audio modalities), audio modality tend to convey more task-specific information than text, such as intonation. However, due to the potential influence of background noise, the quality of these two modalities may varying significantly from different samples. Sun et al. introduces a novel Multimodal Cross and Self-Attention Network (MCSAN)  [138] to dynamically emphasize information from both the linguistic content and acoustic information in speech emotion recognition. Similarily, Hazarika et al. devise a self-attention based multimodal fusion strategy to adapt to the varying quality of textual and audio modalities in emotion recognition task [156]. By weighing modalities at the feature level, the proposed methods provides extra stability against noisy inputs.
@@
Usually applied in vision tasks, channel attention dynamically access the importance of different channels to achieve better performance. To adaptively emphasize important multimodal features while suppressing less relevant ones, Hamid et al. [140] introduces a Multimodal Transfer Module (MMTM) to achieve better overall performance. MMTM is designed for intermediate fusion and can be inserted at different levels of the feature hierarchy, allowing recalibration of channel-wise features of low-quality modality with knowledge from others. A squeeze-then-excitation operation is implemented to generate a joint representation and excitation weights, enabling adaptive feature recalibration.
@@
The motivation of spatial attention is to assess the importance of different locality in the feature map from the channel dimension. Cao et al. introduces a novel lightweight fusion module called Channel Switching and Spatial Attention (CSSA) for multimodal object detection [139]. The proposed CSSA module efficiently fuses inputs from different modalities using spatial attention. In their implementation, spatial attention is enhanced using max and average pooling to retain unique features while effectively fusing information.
@@
As one of the most advanced variants of self-attention, transformers (based on multi-head self-attention) are frequently used in recent multimodal fusion methods to enhance the expressive power of model. Nagrani et al.  proposed Fusion Bottlenecks, a new transformer architecture that restricts cross-modal interactions to bottleneck tokens and mid-late fusion layers only [141]. Fusion Bottlenecks force the model to share only necessary information between modalities, avoiding expensive full pairwise attention. Key to the approach is the insight that full pairwise attention between all modalities’ tokens is often redundant, and condensing interactions through fusion bottlenecks provides an efficient alternative. Girdhar et al. leverages a shared Transformer architecture with shared parameters to handle diverse vision modalities including images, videos, and single-view 3D data. The key idea is to fuse information from different modalities into a shared embedding representation by employing a self-attention mechanism to model cross-modality relationship both temporally and spatially [142]. Wang et al. identify that existing Transformer variants might dilute unimodal modal attention weights on multimodal tasks, leading to sub-optimal overall performance. To overcome this, they propose to dynamically identifies uninformative tokens and replace them with aggregated features from other modalities for effective fusion [143]. Unlike prior works that mainly focus on enhancing reliability of multimodal fusion during training time, Yang et al. proposed to utilizing attention to fully handle corrupted modalities in test time. A confidence-aware loss is devised to enhance credible predictions and mitigate potential influence of noise. The proposed method provides new insights into leveraging attention mechanisms based multimodal fusion for effective test-time adaption.

### v2

@@ 4.2 Quality-discrepancy methods
In addition to the issue of discrepancy in unimodal learning property, even though all modalities convey the same concepts, they contain varying amounts of information related to the target event or objects, resulting a discrepancy in unimodal data quality. Given the inherently greedy nature of deep neural networks, recent studies have further uncovered that multimodal models tend to rely heavily on a single modality with high quality that provides sufficient target-related information, while neglecting the others, leading to their under-optimized [8, 10]. As shown in Fig. 4, Peng et al. [8] empirically verified that the quality of unimodal representations in the multimodal model is worse than that in the corresponding unimodal models, despite the overall superior performance of the multimodal models. Moreover, this under-optimization of each modality is imbalanced, and the modality with high quality is less under-optimized. They claim that this imbalanced problem is caused by the model preference for the high-quality modality with better performance (i.e., the dominant modality). For instance, the audio modality in the curated sound-oriented dataset, VGGSound [130], is less under-optimized compared to the vision modality. Theoretical analysis conducted by Huang et al.  [10] has further confirmed that for the jointly trained multimodal late-fusion framework, the encoder networks can only learn a subset of modalities, and other modalities may fail to be discovered well. This finding is consistent with empirical observations, indicating that the imbalanced preference for different modalities issue needs to be addressed in multimodal learning. To tackle this problem brought by discrepancies in unimodal data quality, a series of studies have been proposed, encompassing learning-objective-based method, optimization-based method, architecture-based method and data-augmentation-based method.
@@ 4.2.1 Learning-objective-based method
Considering the preference of multimodal model for modality with high quality, several methods which introduce additional learning objectives beyond normal joint learning loss function are proposed. These methods tend to utilize additional loss to break the modality preference of multimodal model:
@@
where $\mathcal{L}_{mm}$ is the multimodal loss function and $\mathcal{L}_{uni}$ is a specifically designed loss function that improves unimodal learning, especially the learning of modality with lower quality. Concretely, Yang et al. [117] endowed the multimodal contrastive learning loss to constrain the audio and visual modality to be as close as possible in the semantic space, then help to make full use of all modalities. Also paying attention to the multimodal semantic space, Ma et al[131] proposed a novel regularization technique Calibrating Multimodal Learning, to enforce the consistency between prediction confidence and the number of modalities, avoiding the issue that multimodal classification models tend to over-rely on specific modalities. In addition to these methods in semantic space, some methods introduce the learning objective of knowledge distillation. Du et al. [118] proposed to distill the pre-trained uni-modal features to the corresponding parts of multimodal late-fusion models, to avoid the learning preference for one modality. Differently, Liu et al. [119] designed a self-distillation training strategy, which can automatically teach less-optimized modality from better-optimized modality. Besides these common multimodal classification task, Xu et al [120] focused on the audio-visual fine-grained classification task, which has a higher demand for distinguishable feature distribution. They proposed the multimodal Cosine loss, MMCosine, which performs a modality-wise $L2$ normalization to features and weights towards balanced and better multimodal fine-grained learning. Beyond the generalization performance, Yang et al [121] put attention on the multi-modal robustness. They emphasized that multi-modal models are often vulnerable to attacks on the specific modality, and introduced margin-based regularization to ensure multi-modal robustness.
@@ 4.2.2 Optimization-based method
Optimization-based methods focus on the back-propagation stage. These methods respectively pay attention to the magnitude and direction of unimodal gradient, to improve the learning of modality with lower quality. To slow down the learning of preferred modality with high quality and accordingly offer more training efforts to other modalities, Peng et al. [8] proposed an on-the-fly gradient modulation strategy, which dynamically monitors the contribution difference of various modalities to the final prediction during training and mitigates the gradient magnitude of dominant modality to focus more on other modalities. Inspired by this gradient magnitude modulation idea, Sun et al. [123] extended it into the graph-based models of rumor detection task, adaptively adjusting the magnitude of gradients to alleviate the preference for high-quality modality. Similarly, Fu et al. [124] extended the idea of gradient magnitude modulation into audio-visual video parsing task. They further proposed a modality-separated decision unit to well evaluate the unimodal prediction in the more complex audio-visual video parsing framework with more cross-modal interaction. Besides the gradient magnitude, the direction of unimodal gradient is also explored. Fan et al. [122] suggested that the slow-learning modality’s updating direction is severely disturbed by the preferred high-quality modality, making it hard to its training. Subsequently, they utilize prototypes, which are the centroids of each class within the representation space, to refine the updating direction, aiming to improve unimodal performance.
@@ 4.2.3 Architecture-based method
Architecture-based methods aim to balance the training of modalities with discrepancy in quality by designing better unimodal representation learning modules. He et al. [125] designed a Multimodal Temporal Attention module, which takes into account the temporal influences of all modalities on each individual unimodal branch. This module facilitates the interaction among unimodal branches and achieves adaptive inter-modal balance. Su et al. [126] focused on the generative scenarios, and utilized the coordinated feature space, named Coordinated Knowledge Mining, to improve the under-optimized modalities with the guidance of preferred high-quality modalities. Lin et al. [127] proposed a novel Variational Feature Fusion Module, which regards fusion features as random variables and obtains more balanced segmentation performance for different categories and modalities.
4.2.4 Data-augmentation-based method
Data-augmentation-based methods are with the idea of enhancing low-quality modality at the data input stage. The basic intuition is that solely training of unimodal data could avoid the influence of other modalities. There are two kinds of strategies, introducing additional individual training for less learnt low-quality modality or removing the data of better learnt high-quality modality. For the former idea, Wu et al. [9] firstly measured the relative speed at which the model learns from one modality compared to other modalities, and then introduced the re-balancing steps at which update the underutilized uni-modal branches intentionally in order to accelerate the model to learn from its input modality. For the latter one, Zhou et al. [128] suggested adaptively masking the data input of well-learnt modality, accordingly encouraging the model to fit other modalities better. Most imbalanced multi-modal learning methods ignore the fine-grained modality discrepancy, where the dominant modality could vary across different samples under realistic scenarios. Wei et al. [129] introduced a Shapley-based sample-level modality valuation metric, to observe the contribution of each modality during prediction for each sample, and then dynamically re-sample low-contributing modalities during training. This method achieves fine-grained multi-modal cooperation.
@@ 4.3 Discussion
Overall, balanced multimodal learning methods focus on the discrepancy in learning property or data quality among different modalities, which is caused by the heterogeneity of multimodal data. These methods propose solutions from different perspectives, including learning objective, optimization, architecture, and data augmentation. The balanced multimodal learning is still a sunrise field, with plenty of under-explored directions. For example, it is promising to explore the cooperation between modalities with theoretical guidance. Moreover, the current methods mainly limit to the typical multimodal task, mostly discriminative tasks and a few generative tasks. In fact, beyond these, the Multimodal Large Language Model, which also needs to jointly integrate different modalities, could also suffer from this imbalance problem. It is expected to extend current studies or design new solutions in Multimodal Large Language Model settings.
@@ 5 Dynamic Multimodal Fusion
Current multimodal fusion methods often depends on the assumption that multimodal data are of static quality, which does not always hold in real-world scenarios. Handling multimodal data with dynamically varying quality is an inevitable problem for multimodal intelligence systems. Due to unexpected environmental factors and sensor issues, some modalities may suffer from poor reliability and lack of task-specific information. Besides, the quality of different modalities are dynamically varying from scenarios as shown in Fig. 5. This phenomenon motivates a new multimodal learning paradigm, i.e., dynamical multimodal fusion, which aims to adapt to the dynamically changing quality of multimodal data and selectively integrate task-specific information. In this section, we focus on the challenges of dynamical multimodal fusion and categorize current advances in this literature into three main lines including heuristic, attention-based and uncertainty-aware dynamic fusion.
@@ 5.1 Heuristic dynamic fusion
As a concrete example, illumination conditions can serve as a criterion to introduce dynamic fusion. In around-the-clock applications, previous work identifies that the quality of RGB and thermal modalities are temporally varying in multispectral pedestrian detection task [133]. While the RGB modality tends to contain more useful information compared to the thermal modality under normal illumination conditions, this relationship can be reversed under low light or nighttime conditions where the thermal modality becomes more reliable than RGB. Based on this observation, the authors proposed to utilize an illumination aware fusion module that fuse features from the two modalities adaptively. Similarly, to enable adaptation to varying illumination conditions, Guan et al[132] proposed an illumination-aware weighting mechanism that evaluates the lighting of each input scene. These illumination weights are utilized to dynamically integrate the outputs of sub-networks specialized for day and night detection and segmentation. The proposed method can further be applied to pedestrian detection in security surveillance and autonomous driving systems.
@@
Besides environmental factors like illumination, the criterion of dynamic fusion can be devised according to knowledge about the properties of neural networks. When vision models are involved, activation level of the feature maps and the scaling factors in batch-norm layers can indicate input informativeness. Inspired by this, Li et al[134] proposed an encoder-decoder style network with dynamic fusion strategy. In their implementation, the l1-norm of feature maps is calculated as channel-wise importance, which is then normalized to form a weighting vector that selectively emphasizes informative features from each modality. Wang et al[112] proposed Channel Exchanging Networks (CEN) that dynamically exchange information between modality sub-networks based on channel importance evaluated by the scaling factors of batch-normalize layers. By performing channel exchange within modality-specific regions and sharing convolutional filters, CEN balances unimodal feature learning and multimodal fusion dynamically.
@@
Some relevant works also utilize task-specific modules to achieve dynamic fusion. To highlight a few, Panda et al[135] introduced AdaMML, an adaptive multimodal learning in efficient video recognition. By dynamically selecting optimal input modalities per video segment, this work contributes to data-dependent multimodal selection for efficient video recognition. In this work, a lightweight policy network is devised to dynamically select optimal input modalities for each video segment. AdaMML utilizes Gumbel-Softmax sampling for differentiable training of discrete decisions. Similarly, Xue et al[137] proposed DynMM, a novel approach for dynamic multimodal fusion in deep learning. It introduces modality-level and fusion-level decisions, allowing adaptive selection of input modalities and fusion operations based on data characteristics.
@@ 5.2 Attention-based dynamic fusion
Self-attention models dependency within the input sequence by allowing each element to interact with others. In multimodal emotion recognition (involving textual and audio modalities), audio modality tend to convey more task-specific information than text, such as intonation. However, due to the potential influence of background noise, the quality of these two modalities may vary significantly from different samples. Sun et al. introduced a novel Multimodal Cross and Self-Attention Network (MCSAN)  [138] to dynamically emphasize information from both the linguistic content and acoustic information in speech emotion recognition. Similarly, Hazarika et al. devised a self-attention based multimodal fusion strategy to adapt to the varying quality of textual and audio modalities in emotion recognition task [156]. By weighing modalities at the feature level, the proposed method provides extra stability against noisy inputs.
@@
Usually applied in vision tasks, channel attention dynamically evaluates the importance of different channels to achieve better performance. To adaptively emphasize important multimodal features while suppressing less relevant ones, Hamid et al[140] introduced a Multimodal Transfer Module (MMTM) to achieve better overall performance. MMTM is designed for intermediate fusion and can be inserted at different levels of the feature hierarchy, allowing recalibration of channel-wise features of low-quality modality with knowledge from others. A squeeze-then-excitation operation is implemented to generate a joint representation and excitation weights, enabling adaptive feature recalibration.
@@
The motivation of spatial attention is to assess the importance of different locality in the feature map from the channel dimension. Cao et al. introduced a novel lightweight fusion module called Channel Switching and Spatial Attention (CSSA) for multimodal object detection [139]. The proposed CSSA module efficiently fuses inputs from different modalities using spatial attention. In their implementation, spatial attention is enhanced using max and average pooling to retain unique features while effectively fusing information.
@@
As one of the most advanced variants of self-attention, transformers (based on multi-head self-attention) are frequently used in recent multimodal fusion methods to enhance the expressive power of the model. Nagrani et al.  proposed Fusion Bottlenecks, a new transformer architecture that restricts cross-modal interactions to bottleneck tokens and mid-late fusion layers only [141]. Fusion Bottlenecks force the model to share only necessary information between modalities, avoiding expensive full pairwise attention. Key to the approach is the insight that full pairwise attention between all modalities’ tokens is often redundant, and condensing interactions through fusion bottlenecks provides an efficient alternative. Girdhar et al. leveraged a shared Transformer architecture with shared parameters to handle diverse vision modalities including images, videos, and single-view 3D data. The key idea is to fuse information from different modalities into a shared embedding representation by employing a self-attention mechanism to model cross-modality relationship both temporally and spatially [142]. Wang et al. identified that existing Transformer variants might dilute unimodal modal attention weights on multimodal tasks, leading to sub-optimal overall performance. To overcome this, they propose to dynamically identifies uninformative tokens and replace them with aggregated features from other modalities for effective fusion [143]. Unlike prior works that mainly focus on enhancing reliability of multimodal fusion during training time, Yang et al[yang2023test] proposed to utilizing attention to fully handle corrupted modalities in test time. A confidence-aware loss is devised to enhance credible predictions and mitigate potential influence of noise. The proposed method provides new insights into leveraging attention mechanisms for effective test-time adaption in multimodal fusion.